OpenStudy (anonymous):

Solve 4x^2 = 12x + 40 (Solving Quadratics by Factoring)

OpenStudy (anonymous):

4x^2 - 12x - 40 = 0 x^2 - 3x - 10 = 0 (x-5) (x+2) = 0 x = 5, -2